OpenStudy (anonymous):

A cup of coffee at 183 degrees is poured into a mug and left in a room at 73 degrees. After 8 minutes, the coffee is 136 degrees. Assume that the differential equation describing Newton's Law of Cooling is (in this case) dT/dt=k(T−73). What is the temperature of the coffee after 18 minutes?

OpenStudy (anonymous):

work with the difference in the temperatures. \(183-73=110\) and \(136-73=63\) so your model for the temperature difference it \[T=(\frac{63}{110})^{\frac{t}{8}}\]

OpenStudy (anonymous):

replace \(t\) by 18 to get the difference in the temperature after 18 minutes, then add 73 to get the actual temperature

OpenStudy (anonymous):

you can also use \[T=73+110e^{kt}\] but that takes way more work because you have to solve an exponential equation using logs to find \(k\) . once you truncate \(k\) this method will be less accurate, so don't bother

OpenStudy (anonymous):

Yeah i think i'm supposed to use the second method to get any marks. Just i got stuck, since we don't know the value for dT/dt to find out the value for k.

OpenStudy (anonymous):

ok again work with the difference in the temperatures the model for the difference starting at time 0 will be \(110e^{kt}\) and you solve for \(k\) by replacing \(t\) by 8 and the value by 63 and solving \[63=110e^{8k}\] for k

OpenStudy (anonymous):

you get \[\frac{63}{110}=e^{8k}\] \[8k=\ln(\frac{63}{110}\] \[k=\frac{1}{8}\times \ln(\frac{63}{110})\] and then a calculator

OpenStudy (anonymous):

one you have that \(k\) your model for the difference in the temp will be \[110e^{kt}\] or if you want the actual temp it is \[73+100e^{kt}\] replace \(t\) by 18 to get the answer

OpenStudy (anonymous):

here is k http://www.wolframalpha.com/input/?i=ln%2863%2F110%29%2F8

OpenStudy (anonymous):

ohhh. okay, thanks so much! :)

OpenStudy (anonymous):

..i got the answer to be 111.896 degrees, but it turned out to be wrong. :/

OpenStudy (anonymous):

nvm. i put in 100 instead of 110. >< thanks again though ^^